### v1

# RobotDiffuse: Motion Planning for Redundant Manipulator based on Diffusion Model

Abstract
Redundant manipulators, with their higher Degrees of Freedom (DOFs), offer enhanced kinematic performance and versatility, making them suitable for applications like manufacturing, surgical robotics, and human-robot collaboration. However, motion planning for these manipulators is challenging due to increased DOFs and complex, dynamic environments. While traditional motion planning algorithms struggle with high-dimensional spaces, deep learning-based methods often face instability and inefficiency in complex tasks. This paper introduces RobotDiffuse, a diffusion model-based approach for motion planning in redundant manipulators. By integrating physical constraints with a point cloud encoder and replacing the U-Net structure with an encoder-only transformer, RobotDiffuse improves the model’s ability to capture temporal dependencies and generate smoother, more coherent motion plans. We validate the approach using a complex simulator, and release a new dataset with 35M robot poses and 0.14M obstacle avoidance scenarios. Experimental results demonstrate the effectiveness of RobotDiffuse and the promise of diffusion models for motion planning tasks. The code can be accessed at https://github.com/ACRoboT-buaa/RobotDiffuse.
I INTRODUCTION
Redundant manipulators have garnered significant attention in academia and industry because of their ability to enhance kinematic performance, agility, and dynamic capabilities with higher Degrees of Freedom (DOFs). They can be deployed in various applications [1], including advanced manufacturing, precision assembly, surgical robotics, autonomous material handling, and human-robot collaboration. Motion planning is essential for optimizing the robot’s tasks in real-world settings, ensuring stability and smoothness during operation while considering environmental constraints. Higher DOFs introduce challenges in inverse kinematics during motion planning. Additionally, the complexity and variability of obstacles and the diverse starting and ending points of tasks extend beyond a single, fixed environment.
Motion planning algorithms focus on building a path from the initial to the target position, avoiding environmental obstacles, and satisfying potential constraints while optimizing the path according to certain criteria [2]. Planning algorithms fall mainly into search-based, sampling-based, and intelligence-based. Search-based algorithms can find the theoretical optimal solution but suffer from node explosion in high-dimensional spaces. Sampling-based algorithms use probabilistic completeness to address high-dimensional spaces and have improved sampling strategies to reduce search space and enhance solution efficiency. However, they are still limited by high dimensionality and redundancy, requiring significant computation time for complex problems.
[…]
[width=.7autoplay, loop, controls]10fig/sample/sample133
(a) Sample-based motion planning - RRT*
[…]
[width=.7autoplay, loop, controls]10fig/diffusion/diffusion121
(b) Diffusion-based motion planning - RobotDiffuse
Figure 1: Schematic of the motivation in this paper. Comparing the sample- and our diffusion-based approaches in a dynamic manner. Sample-based ones take more time and backtracking.
Deep learning-based methods learn from experience and make quick decisions by leveraging problem similarities and prior knowledge. Mainstream deep learning motion planning approaches use neural networks to guide sampling directions, iteratively generating the next move until the robot reaches the target. While this simplifies motion planning, as movement is only determined by the target and environment, it has drawbacks: the solving time fluctuates with network modeling capacity and task complexity, and it overlooks previous directions while sampling points, leading to instability and unnecessary wear or inefficiency. Modeling motion planning as a motion generation task, where the neural network generates global movement in one step, offers an effective solution.
The Diffusion Model (DM) is a generative model derived from non-equilibrium thermodynamics proposed in 2015 [21]. In recent years, it has achieved excellent results in video, image, and speech generation, surpassing the Generative Adversarial Network (GAN). The core of the diffusion model is to simulate the evolution of data distribution. During training, it repeatedly adds noise and denoises to learn the transformation process from Gaussian distribution to complex distribution. As shown in Fig. 1, the sample-based methods require extensive exploration of the solution space and backtracking, whereas diffusion models proposed in this paper can quickly generate feasible solutions. The generated results have the advantages of strong generalization and high diversity. However, deploying DM on motion planning faces the following challenges: 1) The U-Net in DM excels at extracting local features but is weak in capturing temporal dependencies. 2) DM ignores physical world constraints and the unique motion characteristics of each robot. Recent attempts [3, 4] employed physical guidance during DM sampling, have not directly addressed either of these challenges, which will be described in related work. This paper proposes a DM-based motion planning approach for redundant manipulators named RobotDiffuse. It describes the physical constraints during training with a well-trained point cloud encoder, accurately capturing obstacles regardless of rotations while filtering out redundant information. Then it uses an encoder-only network to replace the U-Net structure in the DM to learn more temporal dependencies for longer motion planning. We deploy a Pybullet-based simulator [14] and generate a large and complex dataset proposed in this paper. The contribution is listed below.
  • We propose a motion planning approach for redundant manipulators RobotDiffuse, which generates consistent moving directions compared to sample-based methods, and is the first to introduce constraints during training among DM-based methods, as far as we know.
  • Using the point-cloud encoder, RobotDiffuse applies more simplified but complete physical constraints in the training phase.
  • An encoder-only transformer is employed to capture the temporal dependencies in the moving fully, enhancing the coherence of the output planning and improving the model’s ability to understand complex scenes.
  • We release a highly complex motion planning dataset in non-desktop scenarios, with 1.6M frames of poses and 200 obstacle avoidance scenarios.
  • Experimental results demonstrate the effectiveness of our approach and the promise of diffusion models in motion planning tasks.
II Related Work
This section introduces related research on robot motion planning, including those based on search, sampling, and intelligence. In particular, a few methods based on diffusion models are introduced in the intelligent-based part.
Search-based algorithms include A* [5] and D* [6]. The former optimizes search efficiency through a heuristic cost function, sacrificing storage space in exchange, and guides the search process by combining actual distance and estimated distance factors. The latter is effective in dynamic environments but costs higher computational complexity to handle dynamics. Search-based motion planning algorithms can effectively find globally optimal solutions. Still, in high-dimensional spaces, the number of search nodes increases rapidly, leading to the “curse of dimensionality,” where data becomes sparse as dimensions increase, significantly degrading performance.
Sampling-based motion planning algorithms, such as PRM and RRT, are widely used for high-dimensional problems. PRM converts the problem into a graph search by randomly sampling the configuration space and connecting collision-free states. However, the quality of the solution depends on the number of samples, and the method struggles in narrow passages [7, 9]. RRT, while effective for high-dimensional spaces, only guarantees a feasible solution without optimality [8]. To improve this, RRT* introduces rewiring and random reconnection to iteratively refine the path towards asymptotic optimality, though its efficiency is impacted by environmental complexity [10]. Variants like RRT-Connect use bidirectional search and Informed RRT* optimizes sampling within a focused region to avoid unnecessary search [11, 12]. Additionally, strategies like Dai’s direct connection method combine RRT variants to improve path length and clearance [13]. However, these methods face challenges in balancing path quality and computational efficiency, particularly in high-dimensional or time-sensitive scenarios.
Researchers have applied intelligent algorithms to motion planning. Ma used a CGAN-based non-uniform sampling strategy to guide RRT* for 2D planning [16]. While Motion Planning Network [17] and Motion Policy Network [18] have been successful in multi-axis robot planning, their strategies struggle with static environments due to cumulative prediction errors, leading to suboptimal paths and potential planning failures. Recently, diffusion models have attracted attention in the motion planning field due to their excellent performance. [3] samples directly from the posterior trajectory distribution conditioned on task goals, using the diffusion model as a prior. However, it introduced more obstacles but fixed them into a stable position in the dataset. EDMP [4] incorporates a cost function during inference to guide the diffusion process. [15] introduces collision constraint functions into multi-robot motion planning, performed on 2D scenarios. Still, they retain the Unet structure in DM and ignore performing physical constraints during the training phase, which may decline the performance.
III Preliminaries
III-A Problem definition
Notations. The posture state of the robotic arm $x\in\mathbb{R}^{D}$, $D$ is determined by its DOF. A trajectory, $X$, is a sequence of joint states over $N$ frame, i.e. $X\in\mathbb{R}^{D\times N}$. We use $X_{t}$ to represent the noise $T$ sequences of a series of robot joints under $X_{1},X_{2},...X_{T}$. The robot configuration space is defined as $CS\subset\mathbb{R}^{D}$ by the obstacle space $CS_{obs}$ and the accessible space $CS_{free}=CS\backslash CS_{obs}$. The robot’s workspace is denoted as $WS\subset\mathbb{R}^{M}$, where $M$ is the workspace dimension. Similar to configuration space, workspaces can also be divided into obstacle Spaces $WS_{obs}$ and accessibility Spaces $WS_{free}=WS\backslash{WS}_{obs}$. For a motion task, the robot moves in any state $x\in CS_{free}$. The obstacle point cloud is represented as $I\in\mathbb{R}^{K\times 3}$, where $K$ is the number of points in the point cloud. We are interested in solving an optimization problem in the collision-free space $X\subset C_{free}$ after a given task,
$\mathop{min}\limits_{X}J(X;o,CS),$ (1)
where the vector $o$ represents the hyperparameter of the cost function $J$.
Figure 2: RobotDiffuse overview. It contains three steps: embedding, training, and generation. (1) The obstacle point net is embedded into latent space as $z$ by the PointEnet component. Together with init pose $CS_{init}$, target pose $CS_{goal}$ comprises a condition code $C$. (2) $C$ is randomly masked for classifier-free learning and then projected with $t$ into the input token $z_{tk}$. Given a noise and motion $X$, the motion predictor based on an encoder-only transformer predicts the original clean motion $\hat{X}^{1:N}$. (3) Given a condition $C$, we sample random noise $X_{T}$ at the dimensions of the desired motion, then iterate from $T$ to $1$. At each step $t$, the predictor predicts a clean sample $\hat{X}$ and then adds noise making it back to $X_{t-1}$.
IV Method
The framework of our proposed robotDiffuse is shown in Fig. 2. For the task of robot motion generation, the dataset comprises pairs of $(X,C)$, where $C$ is the description of the motion sequence $X$, including obstacle information and the robot’s initial and final pose. Obstacle point cloud environment $I\in{WS}_{obs}$. The insight lies in generating high-quality and diverse robot motion sequences with a desired conditional input $C$, which contains three steps: embedding, training, and sampling. First, the task conditions are embedded in the latent space based on the pre-trained point cloud encoder PointEnet, allowing the model to understand the environmental and task characteristics. Then, Gaussian noises are added to the latent representation, and the process of denoising is performed with a transformer network. The transformer encodes the environment embedding $Z$, initial pose $CS_{init}$, and target pose $CS_{goal}$ with the parameter $\theta^{p}$ to generate a motion trajectory $X$. Finally, the reconstructed latent representations are restored to trajectories and are optimized with our guidance-based approach.
IV-A Condition embedding
To eliminate the interference of irrelevant information such as texture during obstacle avoidance motion planning, we use Compression Autoencoder (CAE) as the point cloud encoder for self-training, which can learn the robust feature space required for planning and control and perform better than other available coding techniques. CAE uses common reconstruction losses and regularization of encoder parameters. To improve the generalization ability, this part of the parameter will be frozen after training, and will not be optimized in subsequent training.
$l_{AE}(\theta^{e},\theta^{d})=\frac{1}{K}\sum_{i\in I}\|i-\hat{i}\|^{2}+\lambda{\sum_{ij}}(\theta_{ij}^{2})^{2},$ (2)
where $\theta^{e}$, $\theta^{d}$ are the parameters of the encoder and decoder, $\lambda$ is the regularization coefficient, and the encoding $\hat{x}$ represents the reconstructed point cloud. PointEnet with 3-layer Perceptrons uses the parameter $\theta^{e}$to compress the obstacle environment information $I\subset WS_{obs}$ into the potential space $Z$, that is,
$Z\leftarrow PointEnet(I;\theta^{e}).$ (3)
IV-B Diffusion Model for Motion Generation
Generative Adversarial Networks (GANs) are known for generating high-quality data through adversarial training, which can be effective when working with limited data. However, they often struggle to generalize under complex conditions. Moreover, predicting robot motion sequences based on tasks is not a simple one-to-one mapping, as multiple motion sequences can correspond to a given task. Inspired by the recent success of diffusion models in image generation, we apply a diffusion model to generate task-conditioned robot joint sequences. This probabilistic framework promotes diverse motion generation, and we leverage the denoising diffusion probabilistic model to facilitate this motion generation process.
RobotDiffuse contains an encoder-only transformer instead of the common Unet structure of diffusion models [19, 20] since it captures more temporal dependencies of motion sequences of any length, as prior work in human motion planning proved [23]. The transformer takes 2 inputs including time step $t$ and condition $C$. Then it predicts a joint sequence from $X_{1},...,X_{T}$. We project the timestep $t$ and condition $C$ into the transformer dimension via a separate feedforward network and integrate them into the token $z_{tk}$. The joint sequence with noise $X_{t}$ is also linearly projected to the transformer size and is embedded and summed with a standard position coding. Input the projected noise to the decoder, by entering $z_{tk}$ across the attention layer, project the result back to the original motion dimension, and as a prediction $\hat{X}_{0}$. Given a task $C$, the pre-trained point cloud Enet encodes $C$ as a potential space $Z$. Now the forward diffusion process can be expressed as follows:
$z_{t}=\sqrt{\overline{\alpha}_{t}}z_{0}+\sqrt{1-\overline{\alpha}_{t}}\epsilon_{t},\epsilon_{t}\sim\mathcal{N}(0,\textbf{I}).$ (4)
$z_{T}$ and the input $X$ are fed into the diffusion model. After the reverse denoising process of step T, the final variable $\hat{z}$ is reconstructed into the robot motion path $\hat{X}$.
We optimize the diffusion model through a two-stage optimization process, combined with motion feedback.
Optimize via Robot Feedback Learning the mean can be reparameterized as learning to predict the original data $x_{0}$. Instead of predicting $\epsilon_{t}$ as formulated by [19], we follow [20] and predict the signal itself. The training loss is defined as a reconstruction loss of $x_{0}$:
$\mathcal{L}_{simple}=\mathbb{E}_{x_{0},n}\|\hat{x}_{\theta}(x_{n},n,c)-x_{0}\|_{1}.$ (5)
As shown in Fig. 2, in denoising step $n$, we concatenate joint pose condition $c$ with joint pose representation $X_{1}^{n},...,X_{T}^{n}$ at noise level $n$, combined with noise embedding as input to a diffusion model, and estimate $x_{0}$.
RobotDiffuse is trained using a composite loss function with 3 components: configuration space loss, geometric task space loss (to force accurate predictions), and collision loss (to prevent catastrophic behavior).
$\mathcal{L}$ $=\lambda_{joint}*\mathcal{L}_{simple}+\lambda_{point}*\mathcal{L}_{point}$ (6)
$+\lambda_{collision}*\mathcal{L}_{collision}.$
$\mathcal{L}_{simple}$ represents the L1 loss of joint rotation, and $\mathcal{L}_{point}$ represents the L1 loss of the robot point cloud, where FK is the Forward Kinematics, which can calculate the position and attitude of any link of the robot given the joint rotation of the robot.
$\begin{gathered}\mathcal{L}_{joint}=\mathcal{L}_{simple}=\frac{1}{N}\sum_{i=1}^{N}\left\|\hat{x}_{0}^{i}-x_{0}^{i}\right\|^{2},\end{gathered}$ (7)
$\mathcal{L}_{point}=\frac{1}{N}\sum_{i=1}^{N}\left\|FK\left(\hat{x}_{0}^{i}\right)-FK\left(x_{0}^{i}\right)\right\|^{2},$ (8)
$\mathcal{L}_{collision}$ represents the collision loss of the point cloud,
$\mathcal{L}_{collision}=\frac{1}{N}\sum_{i=1}^{N}\operatorname{h}\left(FK(\hat{x}_{0}^{i}),I\right).$ (9)
The loss of configuration space has been extensively used in previous work [18]. Geometric task space loss can effectively capture the cumulative error of the motion chain. We construct a distance function $D$. During training, for a given closed surface, $D$ computes the minimum distance between the robot’s point cloud and the surface and returns a negative value if the point cloud is inside the closed surface
$h(ws,I)=\begin{cases}S-D(ws,I)\quad ifD(ws,I)\leq{S}\\ 0\quad ifD(ws,I)>S\end{cases},$ (10)
where $S$ represents the safe distance, and $ws\in WS$ denotes a point corresponding to the robot link, including its position information in the 3-dimensional coordinates.
IV-C Motion Guided Generation
In addition, the motion sequence generated by DDPM is presented in its explicit form, rather than being compressed into the latent space. Given a task, when sampling the model, the noise term $\epsilon_{\theta}(x_{n},n,condition)$ determines the de-noising direction of the motion sequence in each iteration. Therefore, we apply physical constraints on the denoising process to guide the generation process and promote the trajectory to maintain kinematic validity. We combine several different cost functions to guide diffusion. Specifically, at each step of the denoising time, we modify the intermediate trajectory $\tau_{t}$ of the denoiser prediction under the first $t$ diffusion time step. A “noise interpolation” method is proposed to control the initial and final attitude of the robot motion sequence, that is, the prefix frame and suffix frame of $\hat{x}_{0}$ are covered by the initial and final attitude of the task. This encourages motion generation to be consistent with the original input, solves the motion intermediate problem [22], and causes the network to produce a smooth trajectory in the start-goal pair.
V Experiments
V-A Datasets
To verify our approach in more complex obstacle avoidance scenarios, we specifically made a dataset pipeline and produced a dataset Robot-obtalcles-panda (ROP) using Pybullet. ROP consists of more than 35 million frames of robot poses and more than 0.14 million obstacle avoidance scenarios. Each scene sets challenging obstacle avoidance spheres in random positions. This is a highly challenging setup, compared to fixed obstacles in most current work.
This paper proposes the Shared-Tree Informed RRT* algorithm for generating appropriate paths for the dataset, introducing node security defense and shared spanning tree strategies based on Informed RRT*. These innovations improve the performance and quality of traditional motion planning algorithms when applied to redundant manipulators. Each motion planning problem consists of three components: the obstacle scene, initial pose $CS_{init}$, and target end position $WS_{goal}$. Using inverse kinematics, a set of target poses $CS_{goals}$ can be generated. A motion planning algorithm then computes a collision-free joint sequence from the initial pose to the target pose. The shared spanning tree strategy enables simultaneous search of ${CS_{goals}}$, allowing for reuse of the generated tree and ultimately providing an optimal path for multiple goals, thereby further enhancing both generation efficiency and path quality. The node security defense strategy addresses the limitations of traditional algorithms, which often generate paths too close to obstacles, by ensuring the robot avoids collisions and maintains a safe distance from obstacles during movement.
V-B Baselines
The proposed method is compared with several baselines, including the classical sample-based RRT*, intelligence-based approaches, and diffusion-based methods.
RRT* [10] is a random sample-based approach, widely recognized in path planning and known for its ability to provide asymptotically optimal paths in high-dimensional spaces.
MPNet [17] uses autoencoders and multilayer perceptrons for intelligent sampling-based motion planning.
MPINet [18] uses PointNet++ point cloud deep network to model tasks white obstacle point clouds and robot endpoint clouds and utilizes multi-layer perceptron for intelligent motion planning.
MPD [3] is DM-based, pretrains an unconditional motion generator in a fixed multi-obstacle scenario and guides the generation to meet the task requirements through a cost function during sampling.
V-C Evaluation Metrics
In addition to the commonly used success rate, we consider the collision rate and length of the path to evaluate the manipulator’s trajectory in more complete views.
  • Success rate. A trajectory is considered successful if it avoids physical violations and exhibits stable motion, with the final task space posture deviating by no more than 1 cm and 15° from the target. Physical violations include collisions, exceeding joint limits, and erratic behavior (such as excessive twitching).
  • Collision rate. The rate of fatal collisions, which involves both the manipulator itself and surrounding scene obstacles.
  • Path length. For redundant robots, the total motion path is the sum of the individual link trajectories.
V-D Implementation Details
The approach is implemented with Pytorch. The linear layer for condition encoding has input and output dimensions of [4200, 60] and the hidden layer of [786, 512, 256]. We set the potential dimension to 512 and the diffusion step $T$ to 200 to train the diffusion model. We used the Adam optimizer with a learning rate of 0.0003. The number of iterations is 25k. For the transformer module, we used 8 layers with a potential dimension of 512. The training batch size is set to 64. The model is executed on NVIDIA RTX A5000 GPUs.
V-E Results and Analysis
Table I shows a quantitative comparison over 10 minutes on the ROP dataset. RRT* can obtain high-quality solutions with a collision rate of zero within ten minutes, demonstrating its significant potential in redundant robot motion planning and establishing itself as a competitive baseline method. On the other hand, MPINet achieves the highest success rate within the same time frame, indicating the promising prospects of intelligence-based approaches. However, a ten-minute sampling time remains too long for practical applications. In contrast, the diffusion model-based MPD and our RobotDiffuse can achieve high-success-rate shortest paths in just 9 and 15 seconds, underscoring the growing potential of diffusion model-based methods as a promising new approach, as shown in Fig. 3.
TABLE I: Main results on ROP dataset. The arrow indicates that higher or lower values of the metric are preferable.
Success $\uparrow$ (%) Collision $\downarrow$ (%) Length $\downarrow$ Time (s) $\downarrow$
RRT* 73.3 0 15.037 600
MPNet 84.5 14.3 18.492 600
MPINet 87.4 11.6 16.438 600
MPD 79.7 12.4 14.142 9
PointNet[24] 69.4 26.7 17.154 42
DGCNN[25] 73.4 22.5 16.864 64
U-Net[19] 59.3 29.7 17.358 27
Transformer[26] 69.0 16.5 17.137 32
Decoder-only 62.8 23.8 16.195 23
Noise 64.5 41.6 17.628 15
Joint-only 68.1 21.1 16.148 15
Joint&Point 74.0 19.5 14.762 15
JPC 79.1 9.2 14.571 15
Ours 84.9 9.8 14.218 15
Figure 3: Success rate of each method changing over time within ten minutes. Diffusion-based methods can rapidly obtain optimal solutions, whereas RobotDiffuse takes more 6s but gets a better success rate.
Although the success rate of RobotDiffuse may not be significantly higher than that of the baseline methods, the proposed approach stands out for its ability to generate a batch of paths within a limited time frame, regardless of the task or scenario, because of its architectural design. The solution rate reaches as high as 84.9% within 15 seconds, and its success rate surpasses all comparison methods within 5 minutes. After 10 minutes, the success rate is only slightly lower than MPINet, by just 2.5%. Compared to traditional methods, RobotDiffuse shows an 11.6% improvement in success rate.
In terms of path length, our method is 0.819 shorter than the best baseline method based on sampling, demonstrating that motion planning through motion generation can effectively optimize planning results from a global perspective, leading to better planning quality. When compared to MPD, our method is slightly longer by 0.076. Analyzing the output results, MPD produces smoother paths, generally leading to more effective results in most scenarios. However, it fails to make the optimal path in certain situations because its pre-trained large model was developed using a single, unchanging complex scenario. As a result, the model’s output joint motions include unnecessary twists when applied to the simpler scenarios in our dataset.
To summarize, the proposed RobotDiffuse can generate high-quality paths, with shorter path lengths and lower collision rates, which can be more stable in practical applications.
V-F Abation Studies
To validate the effectiveness of our method in terms of point cloud encoders, backbone network of diffusion model, and the introduction of physics guidance, we designed several ablation experiments. All the ablation experiments are trained on ROP datasets and evaluated based on ROP training sets. The results are also shown in Table I.
Effect of PointEnet. In addition to the CAE model, we also tested two other backbones in the step of condition embedding: (1) PointNet model with obstacle point cloud and robot point cloud as inputs; (2) DGCNN model with obstacle point cloud and robot point cloud as inputs. The results indicate that the CAE model is more effective at extracting valuable information from the obstacle point cloud. While PointNet and DGCNN are more mainstream point cloud encoding approaches and perform well on individual tasks [27], they underperform in the tasks presented in this paper. This is likely because these models are supervised and cannot be retrained. In this ablation study, they were jointly trained with the diffusion model, which may have led to information loss during training.
Effect of Diffusion backbones. To validate the effectiveness of incorporating the encoder-only Transformer, we use the traditional U-Net, Transformer, and decoder-only Transformer as predictors in the diffusion model. In the last variant, the condition is introduced through a cross-attention layer. For robotic joint sequence data, the bidirectional attention mechanism in the encoder proves particularly effective, as it enables the model to consider the current joint state concerning both past and future joint states within the sequence. This bidirectional approach allows the encoder to capture long-range dependencies and the global context of joint motion, which is crucial for generating smooth and realistic motion sequences. By integrating information from the entire sequence, the model can more accurately predict the next joint positions or motions. Additionally, relying solely on the encoder further simplifies the overall model structure, reducing computational complexity while enhancing the efficiency and convergence of the modules during training. This streamlined approach contributes to better scalability and performance in real-world applications.
Effect of motion-guide. By comparing the loss function including the $\mathcal{L}_{joint}$, $\mathcal{L}_{joint}+\mathcal{L}_{point}$, and $\mathcal{L}$ without the physical motion guide, this paper demonstrates the effectiveness of each component in the loss function and the guidance. The improvement of RobotDiffuse, based on the predicted sequence itself as proposed in this paper, enhances the success rate in robotic scenarios. Additionally, the incorporation of $\mathcal{L}_{point}$ and $\mathcal{L}_{collision}$ enables the robot to learn attitude error and collision information, which helps the robot generate more effective results. Then, although introducing the diffusion guidance can interfere with RobotDiffuse, leading to an increased collision rate, the overall success rate is higher, primarily because the diffusion guidance addresses challenges in the middle of the motion and allows for better alignment with the target.
[…]
[width=.4autoplay, loop, controls]10fig/movie/movie112
Figure 4: There are two sphere obstacles with a radius of 0.2m in the workspace. The positions of the obstacles are (0, 0.02, 0.63) and (0.57, -0.42, 0.77). Franka Panda moves from the initial pose (1.57, 1.23, 1.68, 1.38, 1.31, 2.85, 1.68) to the target pose (0.21, 1.21, 1.78, 2.45, 1.73, 2.62, 1.52), and ensure the smoothness and safety of the movement, which means that the robot maintains a safe distance of 0.0.5m from the obstacle, and does not overlap with the green area.
V-G Visulaization
For the display of motion planning results, this paper provides a more lightweight visualization scheme, which is based on the Web to set the information of obstacle avoidance tasks and visualize the robot’s motion process, which can provide users with convenient and efficient interaction capabilities and can get started at a low cost, as shown in Fig. 4. This method reads the URDF of the robot based on Three.js and describes the relative motion of the robot’s links based on forward kinematics. For obstacles, it sets the corresponding size of blue geometry to render and uses green geometry to wrap to represent the safe distance.
VI CONCLUSIONS
In this work, we introduced RobotDiffuse, a diffusion-based motion planning approach for redundant manipulators that addresses the challenges of high-dimensional motion planning. By integrating physical constraints through a point cloud encoder and using an encoder-only transformer to capture temporal dependencies, RobotDiffuse enhances both the efficiency and stability of motion planning. Our approach not only overcomes the computational inefficiencies of traditional sampling-based methods but also improves the coherence and quality of the generated trajectories. The proposed method was validated using a complex simulation environment and a large-scale dataset, showing promising results in terms of feasibility and performance. These findings suggest that diffusion models hold significant potential for advancing motion planning tasks, especially in dynamic and complex environments. Future work will focus on further improving the scalability and robustness of the model for real-world applications.